OpenStudy (anonymous):

how would i solve this system of equations algebraically? x^2+y=4 and x+y=2?

OpenStudy (unklerhaukus):

take the difference of the equations, note that it has no y term, solve for x and then substitute back in one of the original equations to solve for y

OpenStudy (anonymous):

stupid of me, got my math final later today, math overload..... ty

OpenStudy (marco26):

get the value of y in terms of x in the equation x+y=2 and substitute it into the second equation x^2+y=4,,you will be able to get x. Substitute in that value of x in any of the two equation and you'll get y

OpenStudy (anonymous):

\[x+y=2\] \[y=2-x\] plug in this in first equation and solve for x :D